OpenStudy (anonymous):

Milla and Luka are 3 kilometers apart and walking toward each other. Milla's average speed is 5 kilometers per hour and Luka's is 4 kilometers per hour. Which equation can be used to find t, the time it takes for Milla and Luka to meet? 5t + 4t = 1 5t + 4t = 3 5t – 4t = 0 5t – 4t = 3

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

Hello @israa88 :) I can see how this problem can be a little confusing, so I will do the best I can to help you out! :) We can start with this, that the 3 kilometers that they start apart with should Not be attached to any letter and Should be in the equation

OpenStudy (anonymous):

so now we are left with b or d

jigglypuff314 (jigglypuff314):

right :) then we have to decide whether those two numbers should be addition or subtraction :/ hmm So think of it this way, if they are walking Towards each other, then they would be closing in on each other Faster (additive)

OpenStudy (anonymous):

so its b :)

OpenStudy (anonymous):

thank u very much

jigglypuff314 (jigglypuff314):

yes, I think so ^_^

jigglypuff314 (jigglypuff314):

Glad I could help! :D

OpenStudy (anonymous):

i have one more question

OpenStudy (anonymous):

can u help me plz

OpenStudy (anonymous):

Jonas jogged up the hill at an average rate of of a mile per minute and then walked down the hill at an average rate of of a mile per minute. The round trip took him 42 minutes. What is the missing value in the table that represents the distance of the trip down the hill? 1/16 x 42/ 16 – x 42 – 1/16 x 1/16 (42 – x)

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

did you copy paste it right? :3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

wait no i am so sorry its the wrong question

OpenStudy (anonymous):

It takes 8 minutes for Byron to fill the kiddie pool in the backyard using only a handheld hose. When his younger sister is impatient, Byron also uses the lawn sprinkler to add water to the pool so it is filled more quickly. If the hose and sprinkler are used together, it takes 5 minutes to fill the pool. Which equation can be used to determine r, the rate in parts per minute, at which the lawn sprinkler would fill the pool if used alone?

OpenStudy (anonymous):

5/8+ 5r = 8 5/8 + 5r = 1 5(2) = r 5/8= 5r

OpenStudy (anonymous):

its a right ?

jigglypuff314 (jigglypuff314):

hmmm I really don't know how to do this one, sorry :(

OpenStudy (anonymous):

its okay thanks for trying though :) @jigglypuff314